Show -5-(+8) on a number line

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Number lines are not algebra.

Answer by ikleyn(53748)

.
-5 - (+8) = -5 -8 = -13.


------------o-----o-----o-----o-----o-----   (Number line)
           -15   -14   -13   -12   -11